# Learning the Universe: Learning to Optimize Cosmic Initial Conditions with Non-Differentiable Structure Formation Models

### v1

@@ 5.2 Reconstructed initial conditions
Figure 5: The quality of the reconstruction is quantified through the power spectra (top) and cross-correlation (bottom) of the reconstructed initial conditions (left), the dark matter overdensity (center), the mass-weighted halo count overdensity and halo count overdensity (right). In the residual panels, we display the transfer function, i.e. the square root of the ratio of the power spectra of the reconstructions with the ground truth. With initial conditions matching to within $10\%$ down to scales of $\sim 0.3h$ Mpc-1, the final density field and the halo fields match down to highly non-linear scales of $\sim 2.1h$ Mpc-1 and $\sim 2.7h$ Mpc-1, respectively. The cross-correlations further demonstrate that high correlation in the initial conditions at large scales ($80\%$ at $\sim 0.23h$ Mpc-1) is sufficient for an $80\%$ correlation in the dark-matter overdensity and halo fields beyond scales of $1h$ Mpc-1. This reflects the gravitational collapse of proto-structures into final structures as discussed in section 5.2.1.
In Fig.4, we visualize the reconstruction after $22$ optimization steps, equivalent to running the full simulation pipeline $96$ times, by which point the data discrepancy has converged (as shown in Appendix C). We show the initial conditions and the corresponding derived fields at redshift $z=0$, including the dark-matter overdensity field, one component of the Lagrangian velocity field (i.e., particles are gridded as their ordering in the initial conditions such that each voxel corresponds to a particle velocity), and the mass-weighted halo count overdensity. To visualize the correlation in the initial conditions we smooth the field over $10h^{-1}$ Mpc with a Gaussian kernel.
@@ 5.2.1 Reconstruction accuracy and temporal mode coupling
[…]
@@ 5.2.2 Peculiar velocities of simulated particles
By reconstructing the initial conditions, we gain access to the complete dynamical evolution of structures, including their velocities at non-linear scales. This is a valuable by-product of our method as peculiar velocities provide a critical tool for testing and differentiating cosmological models, including for example the growth rate of the large-scale structure (see e.g. Koda et al., 2014; Howlett et al., 2017; Boruah et al., 2020). Reconstructions of non-linear peculiar velocities can further enhance the precision of such tests and are also vital for peculiar velocity corrections in local measurements of the expansion rate of the Universe, i.e. the Hubble constant (e.g. Mukherjee et al., 2021; Boruah et al., 2021; Kenworthy et al., 2022; Riess et al., 2022; Peterson et al., 2022). The impact of peculiar velocities on the measurement of $H_{0}$ from gravitational waves and megamasers was studied in Boruah et al. (2021).
In Fig. 6, we show the power spectra, transfer function and cross-correlation of the Eulerian momentum field and the Lagrangian velocity field. For both fields, we average the power spectra and the cross-correlation over the three spatial components of the velocities. Since each particle in the simulation has the same mass, to obtain the momentum field we weigh the particle positions by their velocity when gridding the particles. For the Lagrangian velocity field, we order the particles according to their position in the $3$D grid of the initial conditions and store their velocity in each component.
@@
Figure 7: The halo mass function shows that most of the $M_{200\mathrm{c}}$ halo masses are accurately recovered, with a slight over-estimation in the number of low-intermediate mass haloes and an under-estimation in the intermediate-high mass haloes. The mass resolution limit at $130$ particles, below which the mass is less accurately resolved, is displayed in red. Above the limit, the reconstruction is consistent with $\Lambda$CDM (yellow) within $3\sigma$, as derived from the $1000$ independent simulations in the training data. The residual panel shows the square root of the ratio between the reconstruction and the ground truth, and between the $\Lambda$CDM band and the truth.
@@
Figure 9: The reduced bispectra in two distinct triangular configurations are used to assess the reconstruction accuracy: a configuration defined by fixed wave vector magnitudes as a function of the separation angle $\theta$ (left) and an equilateral triangle configuration with equal wave vector magnitudes to examine all scales (right). The three panels display the results for the final density field, the mass-weighted halo overdensity, and the halo overdensity, along with residuals comparing the reconstructed fields to the ground truth. We also show cosmic variance ($1\sigma$, $2\sigma$, and $3\sigma$ for $\Lambda$CDM) derived from the $1000$ unperturbed, and hence uncorrelated, training simulations (shaded regions). In particular, we note that the largest scales are dominated by cosmic variance. In all six bispectra shown, the reconstructions match the underlying theory and ground truth to within $5\%$ (left) and $10\%$ (right) over most angles and scales.
@@ 5.2.3 Halo mass function
The halo mass distribution, i.e. the number of haloes as a function of mass, serves as an important probe for testing the underlying cosmological model (see, e.g., Jenkins et al., 2001; Artis et al., 2021; Stopyra et al., 2021). Despite our method operating within a non-differentiable structure formation framework, our reconstruction demonstrates accurate recovery across most mass ranges. In Fig. 7, we show the halo mass function of the reconstruction compared with the ground truth and the theoretical $\Lambda$CDM prediction using Tinker (Tinker et al., 2008). We also leverage the $1000$ independent simulations from the training set to estimate the scatter within $\Lambda$CDM in each mass bin. In the study by Mansfield & Avestruz (2021), it was shown that at least $N=130$ particles are required to accurately describe the density profile and mass of $M_{200\mathrm{c}}$ dark matter haloes. For our chosen set of cosmological parameters and number of simulated particles, this corresponds to $8.53\times 10^{13}M_{\odot}h^{-1}$. We show this mass resolution limit with the red shaded region in Fig. 7. While this limit is crucial for recovering accurate density profiles, we note that it is a conservative choice for obtaining an accurate halo mass function. In our case, deviations from the theoretical prediction only begin at $\sim 4\times 10^{13}M_{\odot}h^{-1}$. It is important to note that this arises from the current simulation setup and that our method can also be applied to higher-resolution simulations (see section 6.3).
@@ 5.2.5 Higher-order statistics
where $P$ is the power spectrum, $k_{i}$ are the three wave modes corresponding to the three sides of the triangle, and $B$ is the bispectrum.
@@
Two different triangular configurations typically used (see e.g. Jamieson et al., 2022; Doeser et al., 2024; Jamieson et al., 2024; Bartlett et al., 2025) are chosen to asses the reconstructions. First, we express the bispectrum through the magnitude of two wave vectors, chosen to $k_{1}=0.1h$ Mpc-1 and $k_{2}=1.0h$ Mpc-1 as this captures both large and small scale information, and the separation angle $\theta$, i.e. $Q\left(k_{1},k_{2},\theta\right)$. Second, the bispectrum is expressed through the magnitude of equal wave vectors forming an equilateral triangle, which enables analyzing the fields at all scales. We compute these two forms of the reduced bispectrum for the final over-density field, the mass-weighted halo overdensity and the halo overdensity. Notably, we remove all haloes below the mass resolution limit (see Fig. 7) before computing the bispectra.
@@ 5.3 Application on other ground truths
Up to this point, we have only applied LULO on a single reference simulation with unique initial conditions. In this section, we test the generalizability of our method by applying it to ten other reference simulations. We sample $10$ random white-noise fields and simulate the corresponding new halo catalogues, which are then used to generate the data in the form of the two halo fields.
@@
In Fig. 11, we further show the reconstructed halo mass functions as compared with the respective truth and the $\Lambda$CDM scatter as estimated from the $1000$ training simulations (see section 5.2.3). LULO consistently recovers halo mass functions that agree with the theory, while typically under-predicting the number of low-mass haloes ($M<0.4\times 10^{14}M_{\odot}h^{-1}$; below mass resolution limit) and high-mass haloes ($M>3\times 10^{14}M_{\odot}h^{-1}$) and over-predicting the number of low-intermediate mass haloes. To improve the reconstruction quality, future work will focus on refining the architecture, training strategy, and data likelihoods (see discussion in section 6.2).
Figure 11: Reconstructions of the halo mass function using different ground truth data. The method tends to predict fewer low- and high-mass haloes and a higher number of intermediate-sized haloes than the respective ground truths. In most optimizations and for most masses, the final halo mass function agrees with the underlying theory prediction of $\Lambda$CDM (yellow band) above the mass resolution limit (red band).
@@ 6.1 The importance of modelling non-linear scales
Reconstructions of the initial conditions of our Universe can be leveraged to run constrained simulations that evolve primordial fluctuations to the particular observable large-scale structure distribution of galaxies (see e.g. Wang et al., 2016; Sorce & Tempel, 2018; Gottlöber et al., 2010; Libeskind et al., 2020; Mcalpine et al., 2022; Sawala et al., 2021; Hutt et al., 2022; Ata et al., 2022; Byrohl et al., 2024; Wempe et al., 2024). These simulations enables investigating how structures in our observable Universe formed, thus creating a laboratory to search for model consistencies between $\Lambda$CDM and observations. Recent work includes, for example, analysing the spatial distribution of elliptical galaxies and spiral galaxies (Sawala et al., 2023), intricate properties of clusters (Sorce & Tempel, 2018; Jasche & Lavaux, 2019; Mcalpine et al., 2022; Hutt et al., 2022), fate of clusters (Ata et al., 2022), modified gravity (Bartlett et al., 2021), galaxy intrinsic alignment (Tsaprazi et al., 2022b), large-scale environment of supernovae (Tsaprazi et al., 2022a), gamma-ray emission from dark matter particle annihilation (Kostić et al., 2023a), magnetic fields on cosmological scales (Hutschenreuter et al., 2018), massive neutrinos (Elbers et al., 2023), and the formation of our Local Group (Wempe et al., 2024; Wempe et al., 2025). Re-simulations of our observable Universe also provide access to the complete dynamical evolution of non-linear peculiar velocities, which are essential for peculiar velocity corrections in both Hubble constant measurements using supernova data (e.g., Kenworthy et al., 2022; Riess et al., 2022; Peterson et al., 2022) or using gravitational wave sources (Mukherjee et al., 2021).
@@ 6.3 Flexibility and computational performance of LULO
In this work, to demonstrate the successful integration of non-differentiable models without requiring modifications to the simulators, we used the Gadget-IV code (Springel et al., 2021) and the AHF halo identification algorithm (Gill et al., 2004; Knollmann & Knebe, 2009). The combined computational cost of these with $128^{3}$ simulated particles within a $250h^{-1}$ Mpc box was approximately $5.5$ CPU-hours per simulation, leading to a total of around $27.5$k CPU-hours for generating the training data. The exact number of simulations required for the training data was not explored in this work but remains an area of interest for future investigations. For this setup and having $5000$ training pairs, a single model required approximately $120$ GPU-hours for training.
@@ 7 Conclusions
By fitting explicit physics models that span from the initial conditions to the data, field-level inference surpasses traditional approaches based on statistical summaries while also enabling detailed causal analysis (see e.g. Jasche & Lavaux, 2019; Mcalpine et al., 2022; Wempe et al., 2024). While most field-level approaches rely on differentiable physics simulators, we follow a complementary approach that goes beyond the requirement of differentiability by accelerating the optimization algorithm itself. LULO also separates the neural optimization pipeline from the physics simulator, allowing any simulator to be integrated without extra development.
@@ Acknowledgements
We thank Stuart McAlpine and Matt Ho for useful discussions related to this work and their feedback on the manuscript. This research utilized the Sunrise HPC facility supported by the Technical Division at the Department of Physics, Stockholm University. The computations were also enabled by the Berzelius resource provided by the Knut and Alice Wallenberg Foundation at the National Supercomputer Centre at Linköping University, Sweden. This work has been enabled by support from the research project grant ‘Understanding the Dynamic Universe’ funded by the Knut and Alice Wallenberg Foundation under Dnr KAW 2018.0067. JJ acknowledges the hospitality of the Aspen Center for Physics, which is supported by National Science Foundation grant PHY-1607611. The participation of JJ at the Aspen Center for Physics was supported by the Simons Foundation. JJ acknowledges support by the Swedish Research Council (VR) under the project 2020-05143 – “Deciphering the Dynamics of Cosmic Structure". This work was supported by the Simons Collaboration on “Learning the Universe”.
@@ Appendix B Approximate exact line search
As each function evaluation in the line search can be computationally heavy – as in our case with a full simulation – we aim for a minimal amount of evaluations. We pick an effective line search algorithm called Approximate Exact Line Search (AELS), which is an adaptive variant of golden section search that comes with convergence guarantees (Fridovich-Keil & Recht, 2020). Applying AELS, as seen in Appendix C and Fig. C13 in particular, requires on average $4.3$ simulations per line search.
As part of the development, we also used the bisection bracketing method to benchmark and verify the slightly higher performance of the AELS method. We highlight that AELS works without specifying a search interval and, while being approximate, can be tuned to satisfy accuracy requirements. The hyperparameters that need to be chosen are the initial step size guess $T$, and the adjustment factor $\beta$. We note that having the line search method applied at each optimization step requires setting $T$ at each iteration $t$. We find empirically that $T_{t+1}=\gamma_{t}$ if $\alpha<1$ (see Algorithm 2 in Fridovich-Keil & Recht (2020)) and $T_{t+1}=(\beta^{2}\gamma_{t}+T_{t})/2$ if $\alpha>1$ are effective choices, where $\gamma_{t}$ is the picked step size at iteration $t$. We initialize with $T_{0}=0.01$ and $\beta=0.7$, which slightly differs from the $\beta=2/(1+\sqrt{5})$ chosen in Fridovich-Keil & Recht (2020).
@@
Appendix C Towards optimum
In Fig. C13 we monitor the mean squared error of the mass-weighted and the count-weighted halo fields with their respective ground truths as a function of the number of simulations (lower x-axis) and the number of optimization steps (upper x-axis). The figure displays the optimization process leading to the reconstruction analyzed in seciton 5.2. On average, each line search within one optimization step requires $5$ simulations. Only $4$ optimization steps, corresponding to $20$ simulations, are required to reduce the error, after which further optimization only marginally improves the reconstructions.

### v2

@@ 5.2 Reconstructed initial conditions
[…]
In Fig.4, we visualize the reconstruction after $22$ optimization steps, equivalent to running the full simulation pipeline $96$ times, by which point the data discrepancy has converged (as shown in Appendix C.1). We show the initial conditions and the corresponding derived fields at redshift $z=0$, including the dark-matter overdensity field, one component of the Lagrangian velocity field (i.e., particles are gridded as their ordering in the initial conditions such that each voxel corresponds to a particle velocity), and the mass-weighted halo count overdensity. To visualize the correlation in the initial conditions we smooth the field over $10h^{-1}$ Mpc with a Gaussian kernel.
@@ 5.2.1 Reconstruction accuracy and temporal mode coupling
The transfer function of the initial conditions matches to within $10\%$ for scales $k\leq 0.33h$ Mpc-1, after which it drops. The final density field still matches within $10\%$ to highly non-linear scales of $k=2.09h$ Mpc-1. Similarly, the halo fields stay within $12\%$ for all scales $k\leq 2.69h$ Mpc-1. For comparison, the Nyquist frequency of the box is $k_{N}=2.79h$ Mpc-1, indicating that the agreement extends close to the resolution limit. We note that to explain the data at non-linear scales and obtain $80\%$ correlation in the dark-matter overdensity field, mass-weighted halo overdensity and halo overdensity to scales of $1h$ Mpc-1, $1.28h$ Mpc-1, and $1.15h$ Mpc-1, respectively, the method requires high cross-correlations in the initial conditions at large scales ($80\%$ for $k\leq 0.23h$ Mpc-1). This also connects back to the cross-correlation in Fig. 3 for $\Delta\boldsymbol{x}_{\mathrm{pred}}$ vs $\Delta\boldsymbol{x}_{\mathrm{true}}$, where we identified that linear scales in the initial conditions need to be updated to minimize the data discrepancy at non-linear scales. To highlight the improvement achieved by modeling non-linear scales, we compare our reconstruction to the optimal linear reconstruction in Appendix C.3.
@@ 5.2.2 Peculiar velocities of simulated particles
By reconstructing the initial conditions, we gain access to the complete dynamical evolution of structures, including their velocities at non-linear scales. As Stiskalek et al. (2025) has recently demonstrated, field-level inference of initial conditions using non-linear modelling yields more accurate reconstructions of the peculiar velocity field than alternative methods. Velocities are thus a valuable by-product of our method and provide a critical tool for testing and differentiating cosmological models, including for example the growth rate of the large-scale structure (see e.g. Koda et al., 2014; Howlett et al., 2017; Boruah et al., 2020). Reconstructions of non-linear peculiar velocities can further enhance the precision of such tests and are also vital for peculiar velocity corrections in local measurements of the expansion rate of the Universe, i.e. the Hubble constant (e.g. Mukherjee et al., 2021; Boruah et al., 2021; Kenworthy et al., 2022; Riess et al., 2022; Peterson et al., 2022; Carreres et al., 2024). The impact of peculiar velocities on the measurement of $H_{0}$ from gravitational waves and megamasers was studied in Boruah et al. (2021).
In Fig. 6, we show the power spectra, transfer function and cross-correlation of the Eulerian momentum field and the Lagrangian velocity field. For both fields, we average the power spectra and the cross-correlation over the three spatial components of the velocities. Since each particle in the simulation has the same mass, to obtain the momentum field we weigh the particle positions by their velocity when gridding the particles. For the Lagrangian velocity field, we order the particles according to their position in the 33D grid of the initial conditions and store their velocity in each component.
@@
[…]
@@
[…]
@@ 5.2.3 Halo mass function
[…]
@@ 5.2.5 Higher-order statistics
where $P_{i}\equiv P(k_{i})$ is the power spectrum at $k_{i}$, $k_{i}$ are the three wave modes corresponding to the three sides of the triangle, and $B$ is the bispectrum.
@@
Two different triangular configurations typically used (see e.g. Jamieson et al., 2022; Doeser et al., 2024; Jamieson et al., 2024; Bartlett et al., 2025) are chosen to asses the reconstructions. First, we express the bispectrum through the magnitude of two wave vectors, chosen to $k_{1}=0.1h$ Mpc-1 and $k_{2}=1.0h$ Mpc-1 as this captures both large and small scale information, and the separation angle $\theta$, i.e. $Q\left(k_{1},k_{2},\theta\right)$. Second, the bispectrum is expressed through the magnitude of equal wave vectors forming an equilateral triangle, which enables analyzing the fields at all scales. We compute these two forms of the reduced bispectrum for the final over-density field, the mass-weighted halo overdensity and the halo overdensity. Notably, we remove all haloes below the mass resolution limit (see Fig. 7) before computing the bispectra. In section C.4, we further check the bispectrum obtained for the reconstructed initial conditions.
@@ 5.3 Application on other ground truths
Up to this point, we have only applied LULO on one simulated halo catalogue to reconstruct the underlying cosmic initial conditions. In this section, we test the generalizability of our method by applying it to $10$ other halo catalogues. To generate these, we vary the underlying ground truth initial conditions while keeping the same simulation setup and cosmological parameters. For each new initial conditions, we simulate the corresponding new halo catalogues, which are then used to generate the data in the form of the two halo fields.
@@
[…]
@@ 6.1 The importance of modelling non-linear scales
Reconstructions of the initial conditions of our Universe can be leveraged to run constrained simulations that evolve primordial fluctuations to the particular observable large-scale structure distribution of galaxies (e.g. Lavaux, 2010; Leclercq et al., 2015; Wang et al., 2016; Sorce & Tempel, 2018; Hutt et al., 2022; McAlpine et al., 2022, 2025).
[…]
@@ 6.3 Flexibility and computational performance of LULO
In this work, to demonstrate the successful integration of non-differentiable models without requiring modifications to the simulators, we used the Gadget-IV code (Springel et al., 2021) and the AHF halo identification algorithm (Gill et al., 2004; Knollmann & Knebe, 2009). The combined computational cost of these with $128^{3}$ simulated particles within a $250h^{-1}$ Mpc box was approximately $5.5$ CPU-hours per simulation, leading to a total of around 27.527.5k CPU-hours for generating the training data. The exact number of simulations required for the training data was not explored in this work but remains an area of interest for future investigations. For this setup and having $5000$ training pairs, a single model required approximately $120$ GPU-hours for training.
@@ 7 Conclusions
By fitting explicit physics models that span from the initial conditions to the data, field-level inference surpasses traditional approaches based on statistical summaries while also enabling detailed causal analysis (see e.g. Jasche & Lavaux, 2019; McAlpine et al., 2022; Wempe et al., 2024; McAlpine et al., 2025). While most field-level approaches rely on differentiable physics simulators, we follow a complementary approach that goes beyond the requirement of differentiability by accelerating the optimization algorithm itself. LULO also separates the neural optimization pipeline from the physics simulator, allowing any simulator to be integrated without extra development.
@@ Acknowledgements
We thank Stuart McAlpine, Matt Ho, and Guilhem Lavaux for useful discussions related to this work and their feedback on the manuscript. We also thank Richard Stiskalek and Stephen Stopyra for useful discussions related to the setup of the IC generator and the halo finder algorithm, respectively. This research utilized the Sunrise HPC facility supported by the Technical Division at the Department of Physics, Stockholm University. The computations were also enabled by the Berzelius resource provided by the Knut and Alice Wallenberg Foundation at the National Supercomputer Centre at Linköping University, Sweden. This work has been enabled by support from the research project grant ‘Understanding the Dynamic Universe’ funded by the Knut and Alice Wallenberg Foundation under Dnr KAW 2018.0067. JJ acknowledges the hospitality of the Aspen Center for Physics, which is supported by National Science Foundation grant PHY-1607611. The participation of JJ at the Aspen Center for Physics was supported by the Simons Foundation. JJ acknowledges support by the Swedish Research Council (VR) under the project 2020-05143 – “Deciphering the Dynamics of Cosmic Structure". This work was supported by the Simons Collaboration on “Learning the Universe”.
Data Availability
The data underlying this article will be shared on reasonable request to the corresponding author.
@@ Appendix B Approximate exact line search
As each function evaluation in the line search can be computationally heavy – as in our case with a full simulation – we aim for a minimal amount of evaluations. We pick an effective line search algorithm called Approximate Exact Line Search (AELS), which is an adaptive variant of golden section search that comes with convergence guarantees (Fridovich-Keil & Recht, 2020). Applying AELS, as seen in Appendix C.1 and Fig. C13 in particular, requires on average $4.3$ simulations per line search.
Figure C14: Cross-correlation between the ground truth and the reconstructed initial conditions from our non-linear reconstruction algorithm LULO compared to the optimal linear reconstruction. Using the two halo fields as input, LULO shows a higher cross-correlation than the linear method, here using either the dark-matter field or the mass-weighted halo overdensity as data.
Figure C15: Equilateral bispectrum of the reconstructed initial conditions compared to $1000$ random Gaussian fields and the ground truth. The 11-, 22-, and $3\sigma$ deviations are computed from these random fields and shown in shaded blue. In all $10$ cases (see Section 5.3), the reconstructions show good overall agreement with Gaussian expectations and largely match the respective true initial conditions.
As part of the development, we also used the bisection bracketing method to benchmark and verify the slightly higher performance of the AELS method. We highlight that AELS works without specifying a search interval and, while being approximate, can be tuned to satisfy accuracy requirements. The hyperparameters that need to be chosen are the initial step size guess $T$, and the adjustment factor $\beta$. We note that having the line search method applied at each optimization step requires setting $T$ at each iteration $t$. We find empirically that $T_{t+1}=\gamma_{t}$, where $\gamma_{t}$ is the step size chosen by the algorithm at iteration $t$, if $\alpha<1$ (see Algorithm 2 in Fridovich-Keil & Recht (2020)) and $T_{t+1}=(\beta^{2}\gamma_{t}+T_{t})/2$ if $\alpha>1$ are effective choices. We initialize with $T_{0}=0.01$ and $\beta=0.7$, which slightly differs from the $\beta=2/(1+\sqrt{5})$ chosen in Fridovich-Keil & Recht (2020).
@@
Appendix C Reconstruction performance
In this section, we further examine our reconstruction algorithm LULO. In section C.1, we present the intermediate reconstruction results at each optimization step. In section C.2, we discuss different initialization strategies. To show the high quality of the final reconstruction, we compare our result with the optimal linear reconstruction in section C.3. Finally, in section C.4, we show the bispectrum of the reconstructed initial conditions.
C.1 Towards optimum
In Fig. C13, we monitor the mean squared error of the mass-weighted and the count-weighted halo fields with their respective ground truths as a function of the number of simulations (lower x-axis) and the number of optimization steps (upper x-axis). The figure displays the optimization process leading to the reconstruction analyzed in seciton 5.2. On average, each line search within one optimization step requires $5$ simulations. Only $4$ optimization steps, corresponding to $20$ simulations, are required to reduce the error, after which further optimization only marginally improves the reconstructions.
@@ C.2 Initialization and generalizability
C.3 Comparing to linear reconstruction
To demonstrate the importance of modeling non-linear scales, we compare our reconstruction quality as measured by the cross-correlation with traditional linear methods. Following Legin et al. (2023), we compare our non-linear reconstruction with the optimal linear reconstruction obtained through Wiener filtering. This comparison, shown in Fig. C14, illustrates the improvement achieved by modelling non-linear scales in the data to reconstruct the initial conditions.
A linear reconstruction algorithm assumes that the observed data vector $\boldsymbol{d}$ is related to the true signal $\boldsymbol{x}$ (in our case, the initial conditions) via a linear operator $\mathbf{R}$ and additive noise $\boldsymbol{\varepsilon}$ as
$\boldsymbol{d}=\mathbf{R}\boldsymbol{x}+\boldsymbol{\varepsilon}.$ (20)
In our context, $\mathbf{R}$ corresponds to convolving the initial white-noise field $\boldsymbol{x}$ with the primordial power spectrum, evolving it to redshift $z=0$ using the transfer function, and applying a linear bias to model the halos. Under this linear model and assuming Gaussian statistics, the optimal estimator (see e.g. Zaroubi et al., 1995) is the Wiener filter:
$\mathbf{F}=\langle\boldsymbol{x}\boldsymbol{d}^{\dagger}\rangle\langle\boldsymbol{d}\boldsymbol{d}^{\dagger}\rangle^{-1}=\frac{\mathbf{S}\mathbf{R}}{\mathbf{N}+\mathbf{S}\mathbf{R}^{\top}\mathbf{R}},$ (21)
where $\mathbf{S}$ and $\mathbf{N}$ are the signal and noise covariance matrices. The linear reconstruction is then given by
$\hat{\boldsymbol{x}}_{\mathrm{linear}}=\mathbf{F}\boldsymbol{d},$ (22)
with each element of $\mathbf{F}$ satisfying $\mathbf{F}_{ij}\geq 0$, since covariance matrices are positive semi-definite and the filter cannot amplify signal power.
The cross-correlation between the true initial conditions and the linear reconstruction is thus fundamentally limited by the cross-correlation between the data and the true signal as
$\langle\boldsymbol{x}_{\mathrm{true}}\hat{\boldsymbol{x}}_{\mathrm{linear}}^{\dagger}\rangle=\mathbf{F}\langle\boldsymbol{x}_{\mathrm{true}}\boldsymbol{d}^{\dagger}\rangle\leq\langle\boldsymbol{x}_{\mathrm{true}}\boldsymbol{d}^{\dagger}\rangle.$ (23)
We display this optimal cross-correlation achieved through linear reconstruction in Fig. C14. Our non-linear reconstruction with LULO yields a higher cross-correlation, highlighting the importance of non-linear modelling.
C.4 Bispectrum of reconstructed initial conditions
Our reconstruction, performed without an explicit prior, yields a maximum likelihood estimate of the true initial conditions and faithfully captures the large-scale structure and statistical properties where the data is informative. As shown in the power spectrum in Fig. 5 and Fig. 10, the reconstructed initial conditions exhibit Gaussian random behaviour, characterized by a flat spectrum, down to a certain scale.
Similarly, the reconstructed bispectrum is expected to be consistent with that of a Gaussian white-noise field for certain scales. To assess this, we compare the bispectrum of the reconstructed initial conditions to that of random Gaussian fields. In particular, we compute the equilateral bispectrum for $1000$ random Gaussian fields and show the comparison between these, the ground truth, and the reconstructions in Fig. C15 for the $10$ different ground truths (see section 5.3). Note, we do not display the reduced bispectrum here as the power spectrum of the reconstructed initial conditions goes to zero for smaller scales (see Fig. 5), and the reduced bispectrum would therefore not be informative. We find good overall agreement, with the reconstructed bispectrum closely matching Gaussian expectations, while noting a consistent deviation around k=0.2k=0.2-$0.3h$ Mpc-1. At both larger and smaller scales the bispectrum aligns well with Gaussian predictions and largely agrees with that of the true initial conditions.